# Revisiting kink-like parametrization and constraints using OHD/Pantheon/BAO samples

### v1

@@ Abstract
We reexamine the kink-like parameterization of the deceleration parameter to derive constraints on the transition redshift from cosmic deceleration to acceleration. This is achieved using observational Hubble data, Type Ia supernovae Pantheon samples, and Baryon acoustic oscillations/cosmic microwave background. In this parametrization, the value of the initial $q$ parameter is $q_{i}$, the final value is $q_{f}$, the present value is denoted by $q_{0}$, and the transition duration is given by $\alpha$. We perform our calculations using the Monte Carlo method, utilizing the emcee package. Under the assumption of a flat geometry, we constrain the range of possible values for three scenarios: when $q_{f}$ is unrestricted, when $q_{f}$ is equal to $-1$, and when $\alpha$ is equal to $1/3$. This is done assuming that $q_{i}=1/2$. Here, we achieved that the OHD data fixes the free parameters as in the flat $\Lambda$CDM for unrestricted $q_{f}$. In addition, if we fix $q_{f}=-1$, the model behaves well as the $\Lambda$CDM for the combined dataset. The individual supernova data is causing tension in our determination when contrasted to the $\Lambda$CDM model. We also acquired the current value of the deceleration parameter, which is consistent with the latest results from the Planck Collaboration that assume the $\Lambda$CDM model. Furthermore, we observe a deviation from the standard $\Lambda$CDM model in the current model based on the evolution of $j(z)$, and it is evident that the universe transitions from deceleration to acceleration and will eventually reach the $\Lambda$CDM model in the near future.
@@ I Introduction
We aim to find a parametrization that accurately represents several cosmological models across a broad range of redshift values. In the same vein, the major purpose of this study is to investigate several simple kinematic models for cosmic expansion based on a kink-like parameterization for $q(z)$ [20]. There are some special cases described by our parametrization, such as a flat standard cosmological model with a constant dark energy equation of state ($\omega$CDM), $\Lambda$CDM, the flat DGP braneworld model, the flat quartessence Chaplygin model, modified Polytropic Cardassian model [37].
Our primary focus lies in the transition from cosmic deceleration to acceleration. We are interested in the redshift of this transition, its duration, and the current value of the deceleration parameter. To constrain the parameters of our model, we utilize Observational Hubble data (OHD), Supernovae Pantheon samples, and Baryon Acoustic Oscillations/Cosmic Microwave Background (BAO/CMB). We additionally constrain the value of $H_{0}$ by utilizing datasets rather than assuming uniform or Gaussian priors, as done in previous studies. However, our work is more general and different from similar works in different ways. Firstly, we do not assume the $H_{0}$ as a prior but rather allow our model to behave more generally through observational data. Secondly, in this work, we go one step further by studying the evolution of the equation of state and jerk parameters for a general deceleration parameter. Lastly, we employ the latest Pantheon samples here.
@@ III Observational Data Analysis
Within this section, we employ several combinations of datasets, such as the Observational Hubble data (OHD) data and Type Ia Supernova (SN) distance modulus data, to reconstruct the cosmic deceleration parameter $q$ as a function of the redshift $z$. In addition, we perform a reconstruction of $q$ using the combined $OHD$, $SN$, and $BAO/CMB$ datasets. To explore the parameter space, we will be using the Markov Chain Monte Carlo (MCMC) methodology and Python package emcee [21].
@@ III.1 Observational Hubble Data
III.2 Pantheon SN Ia Sample
We have also opted to work with the Pantheon sample, which represents one of the most extensive collections of SNeIa. This dataset comprises 279 Type Ia supernovae from the PANSTARRS DR1 (PS1) Medium Deep Survey. The SNeIa has redshift values ranging from 0.03 to 0.68. Additionally, distance estimates of SNeIa from the Sloan Digital Sky Survey, SuperNovae Legacy Survey (SNLS), and other low-z and Hubble Space Telescope Samples were combined with PS1 data. Consequently, this combined dataset represents a considerable collection of SNeIa, a total of 1048 supernovae with redshift values ranging from 0.01 to 2.3 [23, 24]. The corresponding chi-square function reads
$\chi^{2}_{SN}=\sum^{N_{SN}}_{i,j=1}\frac{\Delta\mu_{i}}{\sigma_{\mu(z_{i})}},\quad\Delta\mu_{i}=\mu^{th}(\Theta)-\mu_{i}^{obs}.$ (10)
The distance moduli theoretically can be defined as
$\mu^{th}=5\log_{10}D_{L}(z)+\mu_{0},\quad\mu_{0}=5\log_{10}\frac{H_{0}^{-1}}{\mathrm{Mpc}}+25.$ (11)
@@ III.2 Pantheon SN Ia Sample
$D_{L}(z)=\frac{c(1+z)}{H_{0}}S_{K}\bigg{(}H_{0}\int^{z}_{0}\frac{d\overline{z}}{H(\overline{z})}\bigg{)}.$ (12)
The function $S_{K}(x)$ is given by
$S_{K}(x)=\begin{cases}\sinh(x\sqrt{\Omega_{K}})/\Omega_{K},\quad~{}\Omega_{K}>0\\ x,\quad\quad\quad\quad\quad\quad\quad\quad~{}~{}\Omega_{K}=0\\ \sin(x\sqrt{|\Omega_{K}|})/|\Omega_{K}|,~{}~{}\Omega_{K}<0\end{cases}\,.$ (13)
Since we have a spatially flat universe, therefore $\Omega_{K}=0$.
The nuisance parameters in the Tripp formula [25] $\mu=m_{B}-M_{B}+\alpha x_{1}-\beta c+\Delta_{M}+\Delta_{B}$ were recovered using the novel method known as BEAMS with Bias Correction (BBC) [26]. As a result, the observed distance modulus is now equal to the difference between the corrected apparent magnitude $M_{B}$ and the absolute magnitude $m_{B}$ $\left(\mu=m_{B}-M_{B}\right)$. Hence, one can define the chi-square function as follows [27]
$\chi^{2}_{\mathrm{SN}}=\Delta\bm{\mu}^{T}\mathbf{C}^{-1}\Delta\bm{\mu}.$ (14)
The total uncertainty matrix of distance modulus is given by
$\mathbf{C}=\mathbf{D}_{\mathrm{stat}}+\mathbf{C}_{\mathrm{sys}}$ (15)
It is assumed that the diagonal matrix of statistical uncertainties has the following form $\mathbf{D}_{\mathrm{stat},ii}=\sigma^{2}_{\mu(z_{i})}$. Besides, systematic uncertainties are derived using the Bias Corrections (BBC) method, presented in [23].
$\mathbf{C}_{ij,\mathrm{sys}}=\sum^{K}_{k=1}\bigg{(}\frac{\partial\mu^{obs}_{i}}{\partial S_{k}}\bigg{)}\bigg{(}\frac{\partial\mu^{obs}_{j}}{\partial S_{k}}\bigg{)}\sigma^{2}_{S_{k}}.$ (16)
Indexes $\{i,j\}$ denote the redshift bins for distance modulus, $S_{k}$ indicates the magnitude of systematic error, and $\sigma_{S_{k}}$ is its standard deviation uncertainty.
III.3 Baryon Acoustic Oscillations/Cosmic Microwave Background
@@
In this case, the present baryon mass density is denoted by $\Omega_{b0}$, and $\Omega_{\gamma 0}$ is the photon mass density at present. Here, $z_{*}$ is the redshift of photon decoupling, and we use $z_{*}=1091$. To obtain the BAO/CMB constraints, we use the “acoustic scale” $l_{A}=\pi\frac{d_{A}(z_{*})}{r_{s}(z_{*})}$, where $d_{A}(z)=\int^{z}_{0}\frac{dz^{\prime}}{H(z^{\prime})}$ is the angular diameter distance in the comoving coordinates. Here, $D_{V}(z_{BAO})$ is the dilation scale
@@
Finally, we obtain the BAO/CMB constraints $\left(\frac{d_{A}(z_{*})}{D_{V}(z_{BAO})}\right)$ [28]. This dataset was gathered from these references [31, 29, 30].
Consequently, to perform the MCMC sampling, we need to define the chi-square function for our BAO/CMB dataset
$\chi^{2}_{BAO/CMB}=X^{T}C^{-1}X,$ (19)
Where $X$ and $C^{-1}$ are of the form [30]
$X=\left(\begin{array}[]{c}\frac{d_{A}(z_{*})}{D_{V}(0.106)}-30.95\\ \frac{d_{A}(z_{*})}{D_{V}(0.2)}-17.55\\ \frac{d_{A}(z_{*})}{D_{V}(0.35)}-10.11\\ \frac{d_{A}(z_{*})}{D_{V}(0.44)}-8.44\\ \frac{d_{A}(z_{*})}{D_{V}(0.6)}-6.69\\ \frac{d_{A}(z_{*})}{D_{V}(0.73)}-5.45\end{array}\right),\quad C^{-1}=\begin{pmatrix}0.48435&-0.101383&-0.164945&-0.0305703&-0.097874&-0.106738\\ -0.101383&3.2882&-2.45497&-0.0787898&-0.252254&-0.2751\\ -0.164945&-2.454987&9.55916&-0.128187&-0.410404&-0.447574\\ -0.0305703&-0.0787898&-0.128187&2.78728&-2.75632&1.16437\\ -0.097874&-0.252254&-0.410404&-2.75632&14.9245&-7.32441\\ -0.106738&-0.2751&-0.447574&1.16437&-7.32441&14.5022\end{pmatrix}.$ (20)
Additionally, we minimized $\chi^{2}_{OHD}+\chi^{2}_{SN}+\chi^{2}_{BAO}$ to carry out the joint analysis from the combined $OHD+SN+BAO$. The results are numerically derived from MCMC using OHD, Pantheon, BAO, and joint datasets, which are presented in Table 1. Furthermore, the $1-\sigma$ and $2-\sigma$ likelihood contours for the possible subsets of parameter space are presented for different cases in the next section. Here are the priors that we consider for free parameters
@@ III.4 Statistical evaluation
We see that, for $q_{f}$ free case, $\Delta\rm AIC=1.9$ for $OHD$, $2.4$ for $SN$, and $8.0$ for $OHD+SN+BAO$. In this case, the model strongly favors $OHD$ data and is slightly favorable for $SN$. Whereas, in case of $q_{f}=-1$, $\Delta\rm AIC=0.464$ for $OHD$, $1.1$ for $SN$, and $4.6$ for $OHD+SN+BAO$. The model strongly favors all the combinations of datasets. According to BIC analysis, the present model for the case of $q_{f}=-1$ is consistent and favorable.
@@ IV.1 When $q_{f}$ is free
Figure 1: The 1-dimensional and 2-dimensional marginalized posterior distributions in the parameter space $(H_{0},q_{0},z_{t},\alpha)$ for the case when $q_{f}$ is free.
Figure 1: The 1-dimensional and 2-dimensional marginalized posterior distributions in the parameter space $(H_{0},q_{0},z_{t},\alpha)$ for the case when $q_{f}$ is free.
Figure 2: The 1-dimensional and 2-dimensional marginalized posterior distributions in the parameter space $(H_{0},z_{t},\alpha)$ for the case when $q_{f}=-1$.
In FIG. 2, we display 2D marginalized confidence level regions for parameter space $(H_{0},q_{0},z_{t},\alpha)$ using $OHD$, $SN$ and $OHD+SN+BAO$. Finally, we plot the deceleration parameter $q(z)$, the equation of state parameter $\omega(z)$, and the jerk parameter $j(z)$ for all these considered data in FIG. LABEL:fig2. The behavior of $q(z)$ shows a signature flip and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase in the recent past. There is an uninspired phase of the universe when $z\rightarrow-1$ since $q(z)$ diverges more from the de-sitter phase for the combined dataset. Another choice for $q_{f}$ could also be made to ensure the model’s consistency in the distant future. Consequently, an additional analysis is required. The plot also displays the evolution of the equation of state parameter with respect to $z$. It reveals that the $\omega$ is positive at higher redshifts and remains greater than $-1$ (quintessence phase) until the present time for all the datasets but crosses a phantom divide for the far future. The present value of $\omega$ is $-0.72\pm 0.1$, $-0.78^{+0.09}_{-0.07}$, and $-0.76^{+0.08}_{-0.07}$ for $OHD$, $SN$, and $OHD+SN+BAO$, respectively.
However, the function $j(z)=q(2q+1)+(1+z)\frac{dq}{dz}$ demonstrates the deviation of $j$ from the flat $\Lambda$CDM model $(j=1)$ in the best-fit model. The jerk parameter is often used to discriminate against various dark energy models. Observations indicate that, in this model, the current value of $j$ exceeds 1 for all the datasets ($j_{0}=1.3$, $1.71$, $1.7$ for $OHD$, $SN$, $OHD+SN+BAO$, respectively) [36]. Therefore, based on the existing model (where $j\neq 1$ and $q_{0}<0$), it is evident that the dynamic dark energy model considered is the most probable explanation for the current acceleration and needs attention. But in this case, the $j(z)$ is not well constrained as it may be connected to the appearance of abrupt future singularities [34, 35].
@@ IV.2 When $q_{f}$ is $-1$
Let us consider the case in which, besides fixing $q_{i}=1/2$, the final value of the deceleration parameter is fixed to the value $q_{f}=-1$, such that asymptotically in the future, i.e., $z=-1$, the model approaches to a de-sitter phase. Models such as $\Lambda$CDM, DGP, quintessence Chaplygin gas, etc., belong to this class of models [37]. By fixing $q_{f}$, we are left with only three free parameters, namely $(H_{0},z_{t},\alpha)$. We present the $68\%$ and $95\%$ confidence level for this parameter space in FIG. 2. The present values of $q$ for $OHD$, $SN$, and $OHD+SN+BAO$ are obtained explicitly using equation (6), as mentioned in Table 1.
We plot the deceleration parameter $q(z)$, equation of state parameter $\omega(z)$, and jerk parameter $j(z)$ for all these considered data in FIG. LABEL:fig4. The behavior of $q(z)$ shows a signature flip and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase in the recent past and approaches a de-sitter phase as expected. The present values of $\omega$ are obtained as $-0.74^{+0.03}_{-0.06}$, $-0.81^{+0.02}_{-0.04}$, $-0.75^{+0.02}_{-0.03}$ for $OHD$, $SN$, $OHD+SN+BAO$, respectively. It reveals that the $\omega$ is positive at higher redshifts and remains greater than -1 (quintessence phase). It does not cross a phantom divide for the far future in this case.
For this mode, the jerk parameter $j(z)$ is found to be evolving, which indicates a tendency of deviation of the universe from the standard $\Lambda$CDM model. Observations indicate that, in this model, the current value of $j$ exceeds 1 for all the datasets ($j_{0}=1.3$, $1.51$, $1.2$ for $OHD$, $SN$, $OHD+SN+BAO$, respectively). Therefore, based on the existing model (where $j\neq 1$ ), it is evident that the present model seems to converge to 1, i.e., the standard $\Lambda$CDM in the near future. This would indicate that the higher value of $H_{0}$ leads to a lower jerk constraint and decreasing $z_{t}$. The significant deviations from the $\Lambda$CDM predictions suggest a preference for using a parametrization of the dynamical parameter for understanding the dark sector of the universe. It should be noted that the present values of $q_{0}$, $H_{0}$, $z_{t}$ are consistent with the values obtained by many authors [38, 39, 40] (also shown in figures LABEL:fig8 and 5).

### v3

# Revisiting kink-like parametrization and constraints using OHD/Pantheon+/BAO samples

@@ Abstract
We reexamine the kink-like parameterization of the deceleration parameter to derive constraints on the transition redshift from cosmic deceleration to acceleration. This is achieved using observational Hubble data, Type Ia Supernovae Pantheon+ samples and Baryon acoustic oscillations. In this parametrization, the value of the initial $q$ parameter is $q_{i}$, the final value is $q_{f}$, the present value is denoted by $q_{0}$, and the transition duration is given by $\alpha$. We perform our calculations using the Monte Carlo Markov Chain method, utilizing the emcee package. Under the assumption of a flat geometry, we constrain the range of possible values for three scenarios: when $q_{f}$ is unrestricted, when $q_{f}$ is equal to $-1$, and when $\alpha$ is $1/3$. This is done assuming that $q_{i}=1/2$. Here, we achieve that the $SN$ data fixes the free parameters tightly as in the flat $\Lambda$CDM for unrestricted $q_{f}$. In addition, if we fix $q_{f}=-1$, the model behaves well as the $\Lambda$CDM for the combined dataset. We also acquire the current value of the deceleration parameter, which is consistent with the latest results that assume the $\Lambda$CDM model. Furthermore, we observe a deviation from the standard $\Lambda$CDM model in the current model based on the evolution of $j(z)$, and it is evident that the universe transitions from deceleration to acceleration and will eventually reach the $\Lambda$CDM model in the near future.
@@ I Introduction
We aim to find a parametrization that accurately represents several cosmological models across a broad range of redshift values. In the same vein, the major purpose of this study is to investigate several simple kinematic models for cosmic expansion based on a kink-like parameterization for $q(z)$ [20]. There are some special cases described by our parametrization, such as a flat standard cosmological model with a constant dark energy equation of state ($\omega$CDM), $\Lambda$CDM, the flat DGP braneworld model, the flat quartessence Chaplygin model, modified Polytropic Cardassian model [34]. The kink-like parametrization has been investigated using SNeIa: the Gold182 sample processed with the light curve fitter MLCS2k2 and the SNLS survey, analyzed with SALT. There is tension between the results obtained using the GOLD182 dataset and those obtained with the SNLS dataset. It is seen that the SN observations in their work cannot impose strong constraints on the value of $z_{t}$. They remarked that even in the region of interest $(z_{t}\leq 1)$ the difference between the outcomes of the two SNeIa datasets, i.e. SNLS and Gold182, exists and can be related to possible inhomogeneities present in the Gold182 sample. The authors have assumed the old Hubble dataset, which comprises 28 points, and considered the values of $H_{0}$ as uniform priors to constrain the other parameters.
Our primary focus lies in the transition from cosmic deceleration to acceleration. We are interested in the redshift of this transition, its duration, and the current value of the deceleration parameter. To constrain the parameters of our model, we utilize Observational Hubble data (OHD), Supernovae Pantheon+ samples, and Baryon Acoustic Oscillations (BAO). We additionally constrain the value of $H_{0}$ by utilizing datasets rather than assuming uniform or Gaussian priors, as done in previous studies. However, our work is more general and different from similar works in different ways. Firstly, we do not assume the $H_{0}$ as a prior but rather allow our model to behave more generally through observational data. Secondly, in this work, we go one step further by studying the evolution of the equation of state and jerk parameters for a general deceleration parameter. Lastly, we employ the latest Pantheon+ samples here.
@@ III Observational Data Analysis
Within this section, we employ several combinations of datasets, such as the OHD and SN Pantheon+ data, to reconstruct the cosmic deceleration parameter $q$ as a function of the redshift $z$. In addition, we perform a reconstruction of $q$ using the combined $OHD$, $SN$, and $BAO$ datasets. To explore the parameter space, we will be using the MCMC methodology and Python package emcee [21].
@@ III.1 Observational Hubble Data
III.2 Pantheon+ SNeIa Sample
The Pantheon+ dataset comprises distance moduli calculated from 1701 light curves of 1550 SNeIa. These light curves were obtained from 18 different surveys, and they cover a redshift range of $0.001\leq z\leq 2.2613$. It is worth noting that 77 out of the 1701 light curves are associated with Cepheid-containing galaxies [23]. In the context of the Pantheon+ dataset, chi-square values are computed using the equation
$\chi^{2}_{SN}=\Delta\mu^{T}\left(C^{-1}_{stat+sys}\right)\Delta\mu,$ (10)
where $C^{-1}_{stat+sys}$ represents the covariance matrix of the Pantheon dataset, which combines both systematic and statistical uncertainties, $\Delta\mu$ is the distance residual given by $\Delta\mu_{i}=\mu_{i}-\mu_{th}(z_{i})$. It is important to note that $\mu_{i}-m_{i}-M$, where $m_{i}$ is the apparent magnitude and $M$ is the fiducial magnitude of $i^{th}$ SNeIa. The distance moduli theoretically can be defined as
$\mu^{th}=5\log_{10}d_{L}(z)+\mu_{0},\quad\mu_{0}=5\log_{10}\frac{H_{0}^{-1}}{\mathrm{Mpc}}+25,$ (11)
@@ III.2 Pantheon+ SNeIa Sample
$d_{L}(z)=\frac{c(1+z)}{H_{0}}\bigg{(}H_{0}\int^{z}_{0}\frac{d\overline{z}}{H(\overline{z})}\bigg{)}.$ (12)
The Pantheon+ dataset is an improvement over the previous Pantheon sample in that it resolves the ambiguity between the absolute magnitude $M$ and the Hubble constant $H_{0}$. This is accomplished by expressing distance moduli of SNeIa in the Cepheid host, defined as
$\Delta\tilde{\mu}=\begin{cases}\mu_{i}-\mu_{i}^{Ceph}&\text{if }i\in\text{Cepheid hosts}\\ \mu_{i}-\mu_{th}^{z_{i}}&\text{Otherwise}\end{cases}$ (13)
where $\mu_{i}^{Ceph}$ represents Cepheid host of the $i^{th}$ SNeIa which is provided by SH0ES. It is seen that $\mu_{i}-\mu_{i}^{Ceph}$ is sensitive to $M$ and the Hubble constant $H_{0}$. Furthermore, in our analysis, we take $M=-19.523$, which has been determined from SH0ES Cepheid host distances [24].
III.3 Baryon Acoustic Oscillations
@@
In this case, the present baryon mass density is denoted by $\Omega_{b0}$, and $\Omega_{\gamma 0}$ is the photon mass density at present. Here, $z_{*}$ is the redshift of photon decoupling, and we use $z_{*}=1091$. To obtain the BAO constraints, we use the “acoustic scale” $l_{A}=\pi\frac{d_{A}(z_{*})}{r_{s}(z_{*})}$, where $d_{A}(z)=\int^{z}_{0}\frac{dz^{\prime}}{H(z^{\prime})}$ is the angular diameter distance in the comoving coordinates. Here, $D_{V}(z_{BAO})$ is the dilation scale
@@
Finally, we obtain the BAO constraints $\left(\frac{d_{A}(z_{*})}{D_{V}(z_{BAO})}\right)$ [25]. This dataset was gathered from these references [28, 26, 27].
Consequently, to perform the MCMC sampling, we need to define the chi-square function for our BAO dataset
$\chi^{2}_{BAO}=X^{T}C^{-1}X,$ (16)
Where $X$ and $C^{-1}$ are of the form in [27].
Additionally, we minimized $\chi^{2}_{OHD}+\chi^{2}_{SN}+\chi^{2}_{BAO}$ to carry out the joint analysis from the combined $OHD+SN+BAO$. The results are numerically derived from MCMC using Pantheon+, BAO, and joint (OHD+SN+BAO) datasets, which are presented in Table 1. Furthermore, the $1-\sigma$ and $2-\sigma$ likelihood contours for the possible subsets of parameter space are presented for different cases in the next section. Here are the priors that we consider for free parameters
@@ III.4 Statistical evaluation
We see that, for $q_{f}$ free case, $\Delta\rm AIC=0.83$ for $SN$, and $5.7$ for $OHD+SN+BAO$. In this case, the model strongly favours $SN$ data and is a little favourable for the joint samples. Whereas, in case of $q_{f}=-1$, $\Delta\rm AIC=2.56$ for $SN$ and $5.8$ for $OHD+SN+BAO$. In this case, the model favours all the combinations of datasets. According to BIC analysis, the present model for the case of $q_{f}=-1$ is consistent and favourable.
@@ IV.1 When $q_{f}$ is free
Figure 1: The marginalized posterior distributions in the parameter space $(H_{0},q_{0},z_{t},\alpha)$ for the case when $q_{f}$ is free.
Figure 2: The marginalized posterior distributions in the parameter space $(H_{0},z_{t},\alpha)$ for the case when $q_{f}=-1$.
Figure 3: Behavior of the deceleration parameter and the equation of state parameter when $q_{f}$ is free.
In FIG. 2, we display 2D marginalized confidence level regions for parameter space $(H_{0},q_{0},z_{t},\alpha)$ using $SN$ and $OHD+SN+BAO$. Finally, we plot the deceleration parameter $q(z)$, the equation of state parameter $\omega(z)$ for all these considered data in FIG. 3. The behavior of $q(z)$ shows a signature flip, and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase in the recent past. There is an uninspired phase of the universe when $z\rightarrow-1$ since $q(z)$ diverges more from the de-sitter phase. Another choice for $q_{f}$ could also be made to ensure the model’s consistency in the distant future. Consequently, an additional analysis is required. The plot also displays the evolution of the equation of state parameter with respect to $z$. It reveals that the $\omega$ is positive at higher redshifts and remains greater than $-1$ (quintessence phase) until the present time for all the datasets. It does not cross a phantom divide. The present value of $\omega$ is $-0.532^{+0.021}_{-0.018}$ and $-0.54\pm 0.13$, for $SN$ and $OHD+SN+BAO$, respectively.
However, the function $j(z)=q(2q+1)+(1+z)\frac{dq}{dz}$ demonstrates the deviation of $j$ from the flat $\Lambda$CDM model $(j=1)$ in the best-fit model. The jerk parameter is often used to discriminate against various dark energy models. Observations indicate that, in this model, the current value of $j$ is far from $1$ for all the datasets ($j_{0}=0.05$ and $0.02$ for $SN$, $OHD+SN+BAO$, respectively). Therefore, based on the existing model, it is evident that the dynamic dark energy model considered is the most probable explanation for the current acceleration and needs attention. But in this case, the $j(z)$ is not well constrained as it may be connected to the appearance of abrupt future singularities [31, 32].
@@ IV.2 When $q_{f}$ is $-1$
Let us consider the case in which, besides fixing $q_{i}=1/2$, the final value of the deceleration parameter is fixed to the value $q_{f}=-1$, such that asymptotically in the future, i.e., $z=-1$, the model approaches to a de-sitter phase. Models such as $\Lambda$CDM, DGP, quintessence Chaplygin gas, etc., belong to this class of models [34]. By fixing $q_{f}$, we are left with only three free parameters, namely $(H_{0},z_{t},\alpha)$. We present the $68\%$ and $95\%$ confidence levels for this parameter space in FIG. 2. The present values of $q$ for $SN$ and $OHD+SN+BAO$ are obtained explicitly using equation (6), as mentioned in Table 1.
We plot the deceleration parameter $q(z)$, equation of state parameter $\omega(z)$, and jerk parameter $j(z)$ for all these considered data in FIG. 4. The behavior of $q(z)$ shows a signature flip and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase and approaches a de-sitter phase as expected. The present values of $\omega$ are obtained as $-0.56^{+0.007}_{-0.008}$, $-0.58^{+0.005}_{-0.012}$ for $SN$ and $OHD+SN+BAO$, respectively. It reveals that the $\omega$ remains greater than $-1$ (quintessence phase). It does not cross a phantom divide for the far future in this case.
For this model, the jerk parameter $j(z)$ is found to be evolving, which indicates a tendency of deviation of the universe from the standard $\Lambda$CDM model. Observations indicate that, in this model, the current value of $j$ is less than $1$ for all the datasets ($j_{0}=0.5$ and $0.56$ for $SN$, $OHD+SN+BAO$, respectively) [33]. Therefore, based on the existing model (where $j\neq 1$ ), it is evident that the present model seems to converge to $1$, i.e., the standard $\Lambda$CDM in the near future. This would indicate that the higher value of $H_{0}$ leads to a decreased jerk constraint and increased $z_{t}$. The significant deviations from the $\Lambda$CDM predictions suggest a preference for using a parametrization of the dynamical parameter for understanding the dark sector of the universe. It should be noted that the present values of $q_{0}$, $H_{0}$, $z_{t}$ are consistent with the values obtained by many authors [35, 36, 37] (also shown in FIGs. 8).
Figure 4: Behavior of the deceleration parameter, the equation of state parameter and the jerk parameter for the case of $q_{f}=-1$. The upper panel shows the behavior for $SN$, and the lower panel is for $OHD+SN+BAO$.